OpenStudy (anonymous):

**GOLD MEDAL WILL BE GIVEN TO WHO HELPS :) ** Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 A) -4x - 9y = -19 -10y = -30 B) 4x + 3y = 5 2y = -6 C) 7x - 3y = -11 9x = -6 D) 12x - 3y = -33 14x = -28

OpenStudy (anonymous):

@manutd30

OpenStudy (danjs):

multiply the first equation by 3 , then add them together...

OpenStudy (danjs):

The basic rules are, you can multiply an equation by a constant value rearrange equations and add equations together...

OpenStudy (anonymous):

@DanJS what do you mean multiply them?

OpenStudy (danjs):

If you multiply an entire equation by a constant number, the equation is unchanged... \[3*[4x-y=-11]\]

OpenStudy (danjs):

So when you do that then add the two together, a variable goes to zero and drops out

OpenStudy (danjs):

The result after multiplying the first by 3 is there, and now add the two together... 12x - 3y = -33 2x + 3y = 5 ------------------

OpenStudy (danjs):

14x = -29

OpenStudy (danjs):

you want to find a number you can multiply each equation by, so that when added together, you have a variable drop out. In this case -3y + 3y = 0y

OpenStudy (anonymous):

@DanJS so I would have to do this to each one ?

OpenStudy (danjs):

Here i will do each basic step for this example.....

OpenStudy (danjs):

4x - 1y = -11 2x + 3y = 5 If added together, you just will get 6x + 2y = -6

OpenStudy (danjs):

you see that so far?

OpenStudy (anonymous):

yes I see @DanJS

OpenStudy (danjs):

ok, so think of some number you can multiply one of the equations by , so that when added will result in zero for one variable...

OpenStudy (danjs):

if you multiply the first by 3, you will have a -3y in the first and a +3y in the second

OpenStudy (danjs):

you see what i mean ?

OpenStudy (danjs):

3*[4x - y = -11] 1*[2x + 3y = 5] ------------------- that is the same as the original, you can multiply both sides of the equation by a number

OpenStudy (anonymous):

@DanJS yes I see! so it would leave me with Answer B correct??

OpenStudy (danjs):

no, lets continue....

OpenStudy (danjs):

If you carry out what is above.. you get 12x - 3y = -11 2x + 3y = 5 ---------------- That was the point, now when added the y term goes to zero

OpenStudy (danjs):

sorry 12x - 3y = -33 2x + 3y = 5 ------------------

OpenStudy (danjs):

now add them together, both sides of the equals

OpenStudy (anonymous):

14x=28!

OpenStudy (danjs):

right !, remember those two basic rules... and the goal of getting a variable to go to zero like the Y did here

OpenStudy (danjs):

Multiplying an equation by a constant number, and Adding two equations together... that is it

OpenStudy (anonymous):

@DanJS this really helped, thank you for explaining this to me!

OpenStudy (danjs):

I could type all the dumb script rules in, but those are the ideas

OpenStudy (danjs):

One step further.....

OpenStudy (danjs):

Say you have one that has terms on the Y like this ... 3y = 10 ... 2y = 41

OpenStudy (danjs):

How to get those to add and cancel out... multiply the first by +2 Multiply the second by -3

OpenStudy (danjs):

6y -6y ----- 0y

OpenStudy (danjs):

just take care of the X and right side of the = when you multiply through and add too

OpenStudy (danjs):

That is it... good luck

OpenStudy (anonymous):

@DanJS wow you make this very easy, definitely writing these in my notes! thank you so much!

OpenStudy (danjs):

welcome, practice.. hehe